OpenStudy (rulerlol113):

sqrt mn =m(1/2)n(1/2)= can you use fractional exponents to simplify this?

OpenStudy (rulerlol113):

\[\sqrt{mn}=m \frac{ 1 }{ 2 }n \frac{ 1 }{ 2 }=\]

OpenStudy (eli-ite):

yes i think you can use fractional exponents

OpenStudy (rulerlol113):

may i have an example of a similar equation

OpenStudy (baru):

\[\sqrt{mn}=(mn)^\frac{1}{2} =m^ \frac{ 1 }{ 2 }n^ \frac{ 1 }{ 2 }\]

OpenStudy (baru):

there is nothing more you can do with it

OpenStudy (eli-ite):

oh baru got it

OpenStudy (rulerlol113):

OK baru thanks but isnt that a direct answer? dont get in trouble

OpenStudy (baru):

:P

OpenStudy (rulerlol113):

thank you both for your help.i wish i could give two medals

OpenStudy (eli-ite):

your welcome